OpenStudy (anonymous):

4x^2 + 5xy − y^2 = 2 implicit differentiation

OpenStudy (anonymous):

8x + 5xdy/dx +5y - 2y dy/dx =0 (5x-2y)dy/dx = -8x - 5y dy/dx = ( -8x - 5y) / (5x-2y)

OpenStudy (anonymous):

Differentiate both sides with respect to x. And then re-arrange to solve for dy/dx.\[8x+5y+5x\frac{ dy }{ dx }-2y \frac{ dy }{ dx }=0 \rightarrow 5x\frac{ dy }{ dx }-2y \frac{ dy }{ dx }=-8x-5y\]\[\rightarrow \frac{ dy }{ dx }(5x-2y) =-8x-5y \rightarrow \frac{ dy }{ dx }=-\frac{ 8x+5 }{ 5x-2y }\] @sallyxx

OpenStudy (anonymous):

Sorry, I forgot the y in dy/dx. Should be like this:\[\rightarrow \frac{ dy }{ dx }(5x-2y) =-8x-5y \rightarrow \frac{ dy }{ dx }=-\frac{ 8x+5y }{ 5x-2y }\] @sallyxx

OpenStudy (anonymous):

thank you, sorry to be annoying but then it's asking for the y' of it